OpenStudy (anonymous):

a cylinder of volume V is to be cut from a solid sphere, of radius R show that the maximum value of V is /[4piR^3/3sqrt{3}]/

OpenStudy (anonymous):

\[4piR^3/3\sqrt{3}\]

OpenStudy (anonymous):

did you try to express h (formula for V of cylinder) in a terms of r (cylinder) & R(radius of Sphere)? Take a V' and = 0

OpenStudy (anonymous):

take a derivative of: V=2pi*r^2(R^2-r^2)^1/2 (r is variable) maximize with regards to R

OpenStudy (anonymous):

find a formula for the vol of the cylinder in terms of r and h using the relation R^2 = r^2 + (h/2)^2 where r and h are radius of cylinder base and its height. Differentiate this formula for V wrt h and and equate to 0 to find h corresponding to maxm volume. then substitute for h and r in v= pir^2h This will give the required formula.

OpenStudy (anonymous):

jimmyrep is right - it easier to use present r(radius of cylinder) as: r^2=R^2-(h/2)^2 put in formula for cylinder & take a derivative. V=pi*h*(R^2-h^2/4)